Use the worked example above to help you solve this problem. An airplane has wings, each with area 3.95 m2, designed so that air flows over the top of the wing at 249 m/s and underneath the wing at 225 m/s. Find the mass of the airplane such that the lift on the plane will support its weight, assuming the force from the pressure difference across the wings is directed straight upwards.
Answer #1

Solution) A = 3.95 m^2

V1 = 249 m/s

V2 = 225 m/s

m = ?

From Bernoullis equation

P1 + (1/2)(rho)(V1^2) + (rho)(g)(h1) = P2 + (1/2)(rho)(V2^2) + (rho)(g)(h2)

Here h1 = h2

P2 - P1 = (1/2)(rho)(V1^2 - V2^2)

P2 - P1 = (1/2)(1.225)(249^2 - 225^2)

P2 - P1 = 6967.8 Pa

P = F/A

P2 - P1 = F/A

6967.8 = (mg)/(A)

mg = 6967.8(A) = 6967.8(3.95)

mg = 27522.81

m = (27522.81/9.8)

m = 2808.45 kg